# On the Decidability of Presburger Arithmetic Expanded with Powers

### v1

@@ Abstract
We prove that for any integers $\alpha,\beta>1$, the existential fragment of the first-order theory of the structure $\langle\mathbb{Z};0,1,<,+,\alpha^{\mathbb{N}},\beta^{\mathbb{N}}\rangle$ is decidable (where $\alpha^{\mathbb{N}}$ is the set of positive integer powers of $\alpha$, and likewise for $\beta^{\mathbb{N}}$). On the other hand, we show by way of hardness that decidability of the existential fragment of the theory of $\langle\mathbb{N};0,1,<,+,x\mapsto\alpha^{x},x\mapsto\beta^{x}\rangle$ for any multiplicatively independent $\alpha,\beta>1$ would lead to mathematical breakthroughs regarding base-$\alpha$ and base-$\beta$ expansions of certain transcendental numbers.
@@ 1 Introduction
As noted above, automata-theoretic techniques appear inadequate to establish such statements. We make use instead of mathematical tools from Diophantine approximation and transcendental number theory, in particular Baker’s theorem for linear forms in logarithms, in a manner similar to [4, 20].
@@
To place Theorem 2 in context, consider the case of $\alpha=2$ and $\beta=3$. The constant $\log_{3}(2)$ is a transcendental number that is widely conjectured to be normal (and thus in base $3$, every $-\ell$ pattern of trits should appear within $(A_{n})_{n=0}^{\infty}$ with density $3^{-\ell}$). A fortiori, this would entail that the answer to the first query is always positive. However, normality on its own is not sufficient to settle either of the other two queries.
@@ 2.1 Logical theories
A structure $\mathbb{M}$ consists of a universe $U$, constants $c_{1},\ldots,c_{k}\in U$, predicates $P_{1},\ldots,P_{l}$ where each $P_{i}\subseteq U^{\mu(i)}$ for some $\mu(i)\geq 1$, and functions $f_{1},\ldots,f_{m}$ where each $f_{i}$ has the type $f_{i}\colon U^{\delta(i)}\to U$ for some $\delta(i)\geq 1$. By the language of the structure $\mathbb{M}$ we mean the set of all well-formed first-order formulas constructed from symbols denoting the constants $c_{1},\ldots,c_{k}$, predicates $P_{1},\ldots,P_{l}$, and functions $f_{1},\ldots,f_{m}$, as well as the symbols ,,,,¬,$\forall$,$\exists$,$\land$,$\lor$,$\lnot$,=∀ , ∃ , ∧ , ∨ , ¬ , =. We will additionally write $x\in P$ for a unary predicate $P$ to mean $P(x)$. A term is a well-formed expression constructed from constant, function, and variable symbols. Terms represent elements of the universe. A theory is simply a set of sentences, i.e. formulas without free variables. The theory of the structure $\mathbb{M}$ is the set of all sentences in the language of $\mathbb{M}$ that are true in $\mathbb{M}$. A formula is existential if it is of the form $\exists x_{1},\ldots,x_{m}\colon\varphi(x_{1},\ldots,x_{m})$ for $\varphi$ quantifier-free. The existential fragment of a theory $\mathcal{T}$, which itself is a theory, is the set of all existential formulas belonging to $\mathcal{T}$. Finally, a theory $\mathcal{T}$ is decidable if there exists an algorithm that takes a sentence $\varphi$ and decides whether $\varphi\in\mathcal{T}$.
@@
where $D_{j}\geq 1$ and each $t_{j},h_{j}$ is a $\mathbb{Q}$-linear form, $b_{k}\in\mathbb{Z}$, and ${\sim}_{k}\in\{>,=\}$.
@@ 2.2 Number theory
Let $x\in\mathbb{Z}$ and $p\in\mathbb{N}$ be a prime. Then the $p$-adic valuation of $x$, denoted $\nu_{p}(x)$, is the largest integer $n$ such that $p^{n}$ divides $x$, whereas $p^{n+1}$ does not. By definition, $\nu_{p}(0)=+\infty$. For integers $x,y$ and a prime $p$ we have $\nu_{p}(x+y)\geq\min\{\nu_{p}(x),\nu_{p}(y)\}$. Let $z=\frac{a}{b}\in\mathbb{Q}$ be non-zero with $\gcd(a,b)=1$. Then the (absolute logarithmic) height of $z$ is $h(z)\coloneqq\max\{\log|a|,\log|b|\}$. For $z_{1},\ldots,z_{k}\in\mathbb{Q}_{\neq 0}$ we have that $h(1/z_{i})=h(z_{i})$,
@@ 3 Overview of the results
Given multiplicatively independent $\alpha,\beta\in\mathbb{N}_{>1}$, $z_{1},\ldots,z_{l}\in\{\alpha,\beta\}$, $r,s\geq 0$, $A\in\mathbb{Z}^{r\times l}$, $\mathbf{b}\in\mathbb{Z}^{r}$, $C\in\mathbb{Z}^{s\times l}$, and $\mathbf{d}\in\mathbb{Z}^{s}$, decide whether there exists $\mathbf{z}=(z_{1}^{n_{1}},\ldots,z_{l}^{n_{l}})$ such that $A\mathbf{z}>\mathbf{b}$ and $C\mathbf{z}=d$.
@@
theoremequationsmain Let $\alpha,\beta\in\mathbb{N}_{>1}$ be multiplicatively independent and $z_{1},\ldots,z_{l}\in\{\alpha,\beta\}$ for some $l\geq 1$. Further let $s\geq 1$, $C\in\mathbb{Z}^{s\times l}$, $\mathbf{d}\in\mathbb{Z}^{s}$, and $\mathcal{S}$ be the set of solutions of $C\mathbf{z}=\mathbf{d}$, where $\mathbf{z}=(z_{1}^{n_{1}},\ldots,z_{l}^{n_{l}})$. Then $\mathcal{S}\in\mathfrak{A}$. Moreover, a representation of $\mathcal{S}$ in the form (2) can be effectively computed, with the additional property that $z_{\mu(j)}=z_{\sigma(j)}$ for every $X_{j}$ of the form (3).
@@
$\begin{cases}\vspace*{0.15cm}\frac{h_{i}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})}{z_{2}^{n_{2}}}<z_{1}^{n_{1}}/z_{2}^{n_{2}}-a<\frac{h_{j}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})}{z_{2}^{n_{2}}}\quad\text{for all $(i,j)\in X$}\\ \vspace*{0.15cm}z_{1}^{n_{1}},z_{2}^{n_{2}}>z_{3}^{n_{3}}>\cdots>z_{l}^{n_{l}}\\ \vspace*{0.15cm}h_{i}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})>0\quad\text{for all $i\in I$}\\ n_{2}-n_{3}>N\end{cases}$ (7)
where $X,I$ are finite sets of indices, each $h_{i}$ is a $\mathbb{Q}$-linear form, $a\in\mathbb{Q}_{>0}$, and $N\in\mathbb{N}$. Our algorithm for solving the system (7) proceeds by first inductively solving the sub-system consisting of the inequalities $h_{i}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})<h_{j}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})$ for all $(i,j)\in X$, $z_{3}^{n_{3}}>\cdots>z_{l}^{n_{l}}$, and $h_{i}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})>0$ for $i\in I$. If no such solution exists, then (7) does not have a solution either. Otherwise, let $(m_{3},\ldots,m_{l})$ be a solution to the sub-system. In Section 6, we use arguments from Diophantine approximation to prove that in the latter case the system  (7) does always have a solution, and show to construct such a solution from $(m_{3},\ldots,m_{l})$.
In Section 7 we prove Theorem 2 and give a whole class of queries that become decidable assuming decidability of the existential fragment of $\mathcal{PA}(\alpha^{x},\beta^{x})$. These include occurrence of a given pattern in base-$\beta$ or base-$\alpha$ expansions of $\log_{\alpha}(\beta)$. Finally, in Section 8 we show that for any multiplicatively independent $\alpha,\beta\in\mathbb{N}_{>1}$, already for formulas with three alternating blocks of quantifiers (i.e. formulas with quantifier alternation depth 2) the membership problem in $\mathcal{PA}(\alpha^{\mathbb{N}},\beta^{\mathbb{N}})$ is undecidable. This result is included for the sake of describing the decidability landscape as completely as possible. We use the approach developed by Hieronymi and Schulz in [12], but reduce from the Halting Problem for 2-counter machines as opposed to the Halting Problem for Turing machines, which results is a simpler construction. Thus for any multiplicatively independent $\alpha,\beta$, the decidability question for $\mathcal{PA}(\alpha^{\mathbb{N}},\beta^{\mathbb{N}})$ remains open only for formulas containing exactly two alternating blocks of quantifiers.
@@ 6 Adding inequalities
Choose $\mu=\beta^{m_{1}}$ and $\delta\in\mathbb{Q}$ such that $0<\delta<\nu\beta^{m_{1}}$. It remains to argue the correctness of our choice of $\mu,\delta$.
@@
$\biggl{|}\frac{\alpha^{m_{\alpha}}}{\beta^{m_{\beta}}}-1\biggr{|}=\frac{1}{\beta^{m}}\biggl{|}\frac{\alpha^{k}}{\beta^{n_{1}}}-\mu\biggr{|}<\nu.$
@@
Let $\mathbf{z}=(z_{1}^{m_{1}},\ldots,z_{l}^{m_{l}})$ be as above. For $1\leq j\leq r$, define the form
@@
We proceed by induction on $l$. If $l=1$, then the result is immediate. Suppose $l\geq 2$. Write the system $A\mathbf{z}=b\>\land\>C\mathbf{z}>d$ in the form
@@
7 Hardness results for $\mathcal{PA}(\alpha^{x},\beta^{x})$
@@
By symmetry, it is sufficient to prove the proposition for $(A_{n})_{n=0}^{\infty}$. Fix $k\geq 1$, and let $w\in\{0,\ldots,\alpha-1\}^{k}$. Denote by $\lambda(w)$ be the natural number whose base-$\alpha$ expansion equals $w$, and let $b=\lfloor\log_{\beta}(\alpha)\rfloor$. That $w$ occurs at position $n$ in $(A_{n})_{n=0}^{\infty}$ can be expressed as
$\lambda(w)\leq\alpha^{n+k}\cdot\{\log_{\beta}(\alpha)\}<\lambda(w)+1$
@@
$\lambda(w)\leq\alpha^{n+k}(\log_{\beta}(\alpha)-b)<\lambda(w)+1$ $\Leftrightarrow\lambda(w)+b\alpha^{n+k}\leq\alpha^{n+k}\log_{\beta}(\alpha)<\lambda(w)+b\alpha^{n+k}+1$
$\Leftrightarrow\beta^{\lambda(w)}\cdot\beta^{b\alpha^{n+k}}\leq\alpha^{\alpha^{n+k}}<\beta^{\lambda(w)}\cdot\beta^{1+b\alpha^{n+k}}.$
Recall that for any constant $c$ and a term $t$, we can express $c\cdot t$ as $\underbrace{t+\cdots+t}_{c\textrm{ times}}$. Hence the formula
$\varphi(n,y_{0},\ldots,y_{k-1})\coloneqq\bigvee_{\begin{subarray}{c}w=w_{0}\cdots w_{k-1}\\ w\in\{0,\ldots,\alpha-1\}^{k}\end{subarray}}\>$ $\bigwedge_{i=0}^{k-1}y_{i}=w_{i}\>\>\land\>\>$
$\beta^{\lambda(w)}\cdot\beta^{b\alpha^{n+k}}\leq\alpha^{\alpha^{n+k}}<\beta^{\lambda(w)}\cdot\beta^{1+b\alpha^{n+k}}$
defines $(A_{n})_{n=0}^{\infty}$ as required. ∎
@@ 8 Undecidability of $\mathcal{PA}(\alpha^{\mathbb{N}},\beta^{\mathbb{N}})$
A 2-counter machine Minsky machine is given by $R>0$ instructions, numbered $1,\ldots,R$, and two counters $c^{(1)},c^{(2)}$ that take values in $\mathbb{N}$. Each instruction except the $R$th one is either of the form $c^{(i)}=c^{(i)}+1;\>\mathsf{GOTO}\>r$, or of the form $\mathsf{IF}\>c^{(i)}=0\>\mathsf{GOTO}\>r\>\mathsf{ELSE\>c^{(i)}=c_{i}^{(i)}-1;\>GOTO}\>\widetilde{r}$ where $i\in\{1,2\}$ and $r,\widetilde{r}\in\{1,\ldots,R\}$. The execution starts at line $r=1$ with both counters set to zero, and halts if line $r=R$ is reached. Denote by $c_{i}^{(n)}$ the value of the counter $c_{i}$ and by $r_{n}$ the current instruction number after $n$ steps. We refer to $(c^{(1)}_{n},c^{(2)}_{n},r_{n})$ as the configuration of the machine at time $n$. The transition function $f\colon\mathbb{N}\times\mathbb{N}\times\{1,\ldots,R\}\to\mathbb{N}\times\mathbb{N}\times\{1,\ldots,R\}$ of the machine describes how the configuration is updated. By definition, we have that $c^{(1)}_{0}=c^{(2)}_{0}=0$ and $r_{0}=1$.

### v2

@@ Abstract
We prove that for any integers $\alpha,\beta>1$, the existential fragment of the first-order theory of the structure $\langle\mathbb{Z};0,1,<,+,\alpha^{\mathbb{N}},\beta^{\mathbb{N}}\rangle$ is decidable (where $\alpha^{\mathbb{N}}$ is the set of positive integer powers of $\alpha$, and likewise for $\beta^{\mathbb{N}}$). On the other hand, we show by way of hardness that decidability of the existential fragment of the theory of $\langle\mathbb{N};0,1,<,+,x\mapsto\alpha^{x},x\mapsto\beta^{x}\rangle$ for any multiplicatively independent $\alpha,\beta>1$ would lead to mathematical breakthroughs regarding base-$\alpha$ and base-$\beta$ expansions of certain transcendental numbers. Finally, modifying the original proof of Hieronymi and Schulz we show that for any multiplicatively independent $\alpha,\beta>1$, it is undecidable whether a given formula with at most 3 alternating blocks of quantifiers holds in $\langle\mathbb{N};0,1,<,+,\alpha^{\mathbb{N}},\beta^{\mathbb{N}}\rangle$.
@@ 1 Introduction
As noted above, automata-theoretic techniques appear inadequate to establish such statements. We make use instead of mathematical tools from Diophantine approximation and transcendental number theory, in particular Baker’s theorem on linear forms in logarithms, in a manner similar to [4, 20].
@@
To place Theorem 2 in context, consider the case of $\alpha=2$ and $\beta=3$. The constant $\log_{3}(2)$ is a transcendental number that is widely conjectured to be normal (and thus in base $3$, every length-$l$ pattern should appear within $(A_{n})_{n=0}^{\infty}$ with density $3^{-l}$). A fortiori, this would entail that the answer to the first query is always positive. However, normality on its own is not sufficient to settle either of the other two queries.
@@ 2.1 Logical theories
A structure $\mathbb{M}$ consists of a universe $U$, constants $c_{1},\ldots,c_{k}\in U$, predicates $P_{1},\ldots,P_{l}$ where each $P_{i}\subseteq U^{\mu(i)}$ for some $\mu(i)\geq 1$, and functions $f_{1},\ldots,f_{m}$ where each $f_{i}$ has the type $f_{i}\colon U^{\delta(i)}\to U$ for some $\delta(i)\geq 1$. By the language of the structure $\mathbb{M}$ we mean the set of all well-formed first-order formulas constructed from symbols denoting the constants $c_{1},\ldots,c_{k}$, predicates $P_{1},\ldots,P_{l}$, and functions $f_{1},\ldots,f_{m}$, as well as the symbols ,,,,¬,$\forall$,$\exists$,$\land$,$\lor$,$\lnot$,=∀ , ∃ , ∧ , ∨ , ¬ , =. We will additionally write $x\in P$ for a unary predicate $P$ to mean $P(x)$. A term is a well-formed expression constructed from constant, function, and variable symbols. Terms represent elements of the universe. A theory is simply a set of sentences, i.e. formulas without free variables. The theory of the structure $\mathbb{M}$ is the set of all sentences in the language of $\mathbb{M}$ that are true in $\mathbb{M}$. A formula is existential if it is of the form $\exists x_{1}\cdots\exists x_{m}\colon\varphi(x_{1},\ldots,x_{m})$ for $\varphi$ quantifier-free. The existential fragment of a theory $\mathcal{T}$, which itself is a theory, is the set of all existential formulas belonging to $\mathcal{T}$. Finally, a theory $\mathcal{T}$ is decidable if there exists an algorithm that takes a sentence $\varphi$ and decides whether $\varphi\in\mathcal{T}$.
@@
where $D_{j}\geq 1$ and each $t_{j},h_{j}$ is a $\mathbb{Q}$-linear form, $c_{k}\in\mathbb{Z}$, and ${\sim}_{k}\in\{>,=\}$.
@@ 2.2 Number theory
Let $x\in\mathbb{Z}$ and $p\in\mathbb{N}$ be a prime. Then the $p$-adic valuation of $x$, denoted $\nu_{p}(x)$, is the largest integer $n$ such that $p^{n}$ divides $x$, whereas $p^{n+1}$ does not. By convention, $\nu_{p}(0)=+\infty$. For integers $x,y$ and a prime $p$ we have $\nu_{p}(x+y)\geq\min\{\nu_{p}(x),\nu_{p}(y)\}$. Let $z=\frac{a}{b}\in\mathbb{Q}$ be non-zero with $\gcd(a,b)=1$. Then the (absolute logarithmic) height of $z$ is $h(z)\coloneqq\max\{\log|a|,\log|b|\}$. For $z_{1},\ldots,z_{k}\in\mathbb{Q}_{\neq 0}$ we have that $h(1/z_{i})=h(z_{i})$,
@@ 3 Overview of the results
Given multiplicatively independent $\alpha,\beta\in\mathbb{N}_{>1}$, $z_{1},\ldots,z_{l}\in\{\alpha,\beta\}$, $r,s\geq 0$, $A\in\mathbb{Z}^{r\times l}$, $\mathbf{b}\in\mathbb{Z}^{r}$, $C\in\mathbb{Z}^{s\times l}$, and $\mathbf{d}\in\mathbb{Z}^{s}$, decide whether there exists $\mathbf{z}=(z_{1}^{n_{1}},\ldots,z_{l}^{n_{l}})$ such that $A\mathbf{z}>\mathbf{b}$ and $C\mathbf{z}=\mathbf{d}$.
@@
theoremequationsmain Let $\alpha,\beta\in\mathbb{N}_{>1}$ be multiplicatively independent and $z_{1},\ldots,z_{l}\in\{\alpha,\beta\}$ for some $l\geq 1$. Further let $s\geq 1$, $C\in\mathbb{Z}^{s\times l}$, $\mathbf{d}\in\mathbb{Z}^{s}$, and $\mathcal{S}\subseteq\mathbb{N}^{l}$ be the set of solutions of $C\mathbf{z}=\mathbf{d}$, where $\mathbf{z}=(z_{1}^{n_{1}},\ldots,z_{l}^{n_{l}})$. Then $\mathcal{S}\in\mathfrak{A}$. Moreover, a representation of $\mathcal{S}$ in the form (2) can be effectively computed, with the additional property that $z_{\mu(j)}=z_{\sigma(j)}$ for every $X_{j}$ of the form (3).
@@
$\begin{cases}\vspace*{0.15cm}\frac{h_{i}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})}{z_{2}^{n_{2}}}<z_{1}^{n_{1}}/z_{2}^{n_{2}}-a<\frac{h_{j}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})}{z_{2}^{n_{2}}}\quad\text{for all $(i,j)\in X_{1}\times X_{2}$}\\ \vspace*{0.15cm}z_{1}^{n_{1}},z_{2}^{n_{2}}>z_{3}^{n_{3}}>\cdots>z_{l}^{n_{l}}\\ \vspace*{0.15cm}h_{i}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})>0\quad\text{for all $i\in I$}\\ n_{2}-n_{3}>N\end{cases}$ (7)
where $X_{1},X_{2},I$ are finite sets of indices, each $h_{i}$ is a $\mathbb{Q}$-linear form, $a\in\mathbb{Q}_{>0}$, and $N\in\mathbb{N}$. Our algorithm for solving the system (7) proceeds by first inductively solving the sub-system consisting of the inequalities $h_{i}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})<h_{j}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})$ for all $(i,j)\in X$, $z_{3}^{n_{3}}>\cdots>z_{l}^{n_{l}}$, and $h_{i}(z_{3}^{n_{3}},\ldots,z_{l}^{n_{l}})>0$ for $i\in I$. If no such solution exists, then (7) does not have a solution either. Otherwise, let $(m_{3},\ldots,m_{l})$ be a solution to the sub-system. In Section 6, we use arguments from Diophantine approximation to prove that in the latter case the system (7) does always have a solution, and show to construct such a solution from $(m_{3},\ldots,m_{l})$.
In Section 7 we prove Theorem 2 and give a whole class of queries that become decidable assuming decidability of the existential fragment of $\mathcal{PA}(\alpha^{x},\beta^{x})$. These include occurrence of a given pattern in base-$\beta$ or base-$\alpha$ expansions of $\log_{\alpha}(\beta)$. Finally, in Section 8 we show that for any multiplicatively independent $\alpha,\beta\in\mathbb{N}_{>1}$, already for formulas with three alternating blocks of quantifiers (i.e. formulas with quantifier alternation depth 2) the membership problem in $\mathcal{PA}(\alpha^{\mathbb{N}},\beta^{\mathbb{N}})$ is undecidable. This result is included for the sake of describing the decidability landscape as completely as possible. We use the approach developed by Hieronymi and Schulz in [12], but reduce from the Halting Problem for 2-counter machines as opposed to the Halting Problem for Turing machines, which results in a simpler construction. Thus for any multiplicatively independent $\alpha,\beta$, the decidability question for $\mathcal{PA}(\alpha^{\mathbb{N}},\beta^{\mathbb{N}})$ remains open only for formulas containing exactly two alternating blocks of quantifiers.
@@ 6 Handling inequalities
Choose $\mu=1/\beta^{m_{1}}$ and $\delta\in\mathbb{Q}$ such that $0<\delta\beta^{m_{1}}<\nu$. It remains to argue the correctness of our choice of $\mu,\delta$.
@@
$\biggl{|}\frac{\alpha^{m_{\alpha}}}{\beta^{m_{\beta}}}-1\biggr{|}={\beta^{m_{1}}}\biggl{|}\frac{\alpha^{m_{\alpha}}}{\beta^{n_{1}}}-\mu\biggr{|}<\nu.$
@@
The “if” direction is trivial as one can take $\mathbf{z}:=\widetilde{\mathbf{z}}$. Thus, we focus on the other direction. Let $\mathbf{z}=(z_{1}^{m_{1}},\ldots,z_{l}^{m_{l}})$ be as above. For $1\leq j\leq r$, define the form
@@
We proceed by induction on $l$. If $l=1$, then the result is immediate. Suppose $l\geq 2$. Write the system $A\mathbf{z}>b\>\land\>C\mathbf{z}=d$ in the form
@@
7 Hardness results for the existential fragment of $\mathcal{PA}(\alpha^{x},\beta^{x})$
@@
By symmetry, it is sufficient to prove the proposition for $(A_{n})_{n=0}^{\infty}$. For $x\geq 1$, denote by $f(x)$ the integer $\alpha^{m}$ such that $\alpha^{m}\leq x<\alpha^{m+1}$, noting that $f(x)=\alpha^{\lfloor x\log_{\alpha}\beta\rfloor}$. Fix $k\geq 1$, and let $w\in\{0,\ldots,\alpha-1\}^{k}$. Denote by $\lambda(w)$ the natural number whose base-$\alpha$ expansion equals $w$. That $w$ occurs at position $n$ in $(A_{n})_{n=0}^{\infty}$ can be expressed as
$\lambda(w)<\{\alpha^{n}\log_{\alpha}\beta\}\cdot\alpha^{k}<\lambda(w)+1$
@@
$\alpha^{\lambda(w)}<\left(\frac{\beta^{\alpha^{n}}}{\alpha^{\lfloor\alpha^{n}\log_{\alpha}\beta\rfloor}}\right)^{\alpha^{k}}<\alpha^{\lambda(w)+1}.$
Recall that
$\alpha^{\lfloor\alpha^{n}\log_{\alpha}\beta\rfloor}=f(\beta^{\alpha^{n}}),$
and for any constant $c$ and a term $t$, we can express $c\cdot t$ as $\underbrace{t+\cdots+t}_{c\textrm{ times}}$. Hence the formulas
$\varphi(n,y_{0},\ldots,y_{k-1})\coloneqq\exists m\colon$ $\alpha^{m}\leq\beta^{\alpha^{n}}<\alpha^{m+1}\>\>\land$
$\alpha^{\lambda(y_{0}\cdots y_{k-1})+m\alpha^{k}}<\beta^{\alpha^{n+k}}<\alpha^{\lambda(y_{0}\cdots y_{k-1})+1+m\alpha^{k}}$
for $k\geq 1$ define $(A_{n})_{n=0}^{\infty}$ as required. ∎
@@ 8 Undecidability of $\mathcal{PA}(\alpha^{\mathbb{N}},\beta^{\mathbb{N}})$
A 2-counter Minsky machine is given by $R>0$ instructions, numbered $1,\ldots,R$, and two counters $c^{(1)},c^{(2)}$ that take values in $\mathbb{N}$. Each instruction except the $R$th one is either of the form $c^{(i)}=c^{(i)}+1;\>\mathsf{GOTO}\>r$, or $\mathsf{IF}\>c^{(i)}=0\>\mathsf{GOTO}\>r\>\mathsf{ELSE\>c^{(i)}=c_{i}^{(i)}-1;\>GOTO}\>\widetilde{r}$ where $i\in\{1,2\}$ and $r,\widetilde{r}\in\{1,\ldots,R\}$. The execution starts at line $r=1$ with both counters set to zero, and halts if the line $r=R$ is reached. Denote by $c_{i}^{(n)}$ the value of the counter $c_{i}$ and by $r_{n}$ the current instruction number after $n$ steps. We refer to $(c^{(1)}_{n},c^{(2)}_{n},r_{n})$ as the configuration of the machine at time $n$. The transition function $f\colon\mathbb{N}\times\mathbb{N}\times\{1,\ldots,R\}\to\mathbb{N}\times\mathbb{N}\times\{1,\ldots,R\}$ of the machine describes how the configuration is updated. By definition, we have that $c^{(1)}_{0}=c^{(2)}_{0}=0$ and $r_{0}=1$.
@@
Acknowledgements.
Toghrul Karimov and Joël Ouaknine were supported by the DFG grant 389792660 as part of TRR 248 (see https://perspicuous-computing.science). Joël Ouaknine is also affiliated with Keble College, Oxford as emmy.network Fellow. James Worrell was supported by EPSRC Fellowship EP/X033813/1. The Max Planck Institute for Software Systems is part of the Saarland Informatics Campus. The authors thank the anonymous referees for their constructive comments, and Mihir Vahanwala and Valérie Berthé for helpful initial discussions.